# RoDE: Linear Rectified Mixture of Diverse Experts for Food Large Multi-Modal Models

### v1

@@ Abstract
Large Multi-modal Models (LMMs) have significantly advanced a variety of vision-language tasks. The scalability and availability of high-quality training data play a pivotal role in the success of LMMs. In the realm of food, while comprehensive food datasets such as Recipe1M offer an abundance of ingredient and recipe information, they often fall short of providing ample data for nutritional analysis. The Recipe1M+ dataset, despite offering a subset for nutritional evaluation, is limited in the scale and accuracy of nutrition information. To bridge this gap, we introduce Uni-Food, a unified food dataset that comprises over 100,000 images with various food labels, including categories, ingredients, recipes, and ingredient-level nutritional information. To mitigate the conflicts arising from multi-task supervision during fine-tuning of LMMs, we introduce a novel Linear Rectification Mixture of Diverse Experts (RoDE) approach. RoDE utilizes a diverse array of experts to address tasks of varying complexity, thereby facilitating the coordination of trainable parameters, i.e., it allocates more parameters for more complex tasks and, conversely, fewer parameters for simpler tasks. RoDE implements linear rectification union to refine the router’s functionality, thereby enhancing the efficiency of sparse task allocation. These design choices endow RoDE with features that ensure GPU memory efficiency and ease of optimization. Our experimental results validate the effectiveness of our proposed approach in addressing the inherent challenges of food-related multitasking.
@@
Figure 2: Label Constitution. The annotation consists of category, ingredient list, recipe instructions, and nutrition information.
Introduction
@@
Food occupies a central position in our daily lives, leading to the emergence of various food-related tasks, e.g., ingredient recognition, recipe generation, and nutritional estimation. These tasks have attracted considerable research interests over the years (Chen and Ngo 2016a; Zhu et al. 2019; Pan et al. 2020; Chen et al. 2020; Min et al. 2023; Luo et al. 2023). Building on the success of Large Language Models (LLMs) (Radford et al. 2018; Touvron et al. 2023), Large Multi-modal Models (Liu et al. 2023a; Luo et al. 2023) have begun to make a significant impact in many specialized areas, including the food domain (Yin et al. 2023).
@@
The success of LLMs and LMMs can be largely attributed to the availability of large-scale training data. In the food domain, while Recipe1M (Salvador et al. 2017) provides over a million web-crawled recipes and thousands of food images, it lacks comprehensive nutritional information. While Recipe1M+ (Marin et al. 2019) includes a subset containing nutritional information, it suffers from limitations in data scale and annotation quality. Similarly, nutrition-focused datasets such as Nutrition5k (Thames et al. 2021) are also limited in scale and may suffer from domain shift issues, as the labeled data primarily consists of lightly processed ingredients rather than fully cooked meals. Moreover, previous studies (Zhou et al. 2022; Chen, Jie, and Ma 2024) have indicated that integrating diverse sources of data for training can lead to data conflicts. The performance of an LMM on a specific task also depends heavily on the representation of that task’s data within the overall training dataset (Chen, Jie, and Ma 2024). These challenges highlight the need for a unified dataset covering all food-related tasks from the same source, aiming at mitigating data conflicts and ensuring balanced representation across tasks.
@@
In order to overcome these obstacles, we introduce Uni-Food, a unified dataset that encapsulates various food-related information, i.e., food category, ingredients, recipes, and valuable ingredient-level nutritional information for each food image. Specifically, we curate 100,000 samples from Recipe1M (Salvador et al. 2017) and employ ChatGPT-4 (Achiam et al. 2023) to generate nutritional information for each ingredient list. This information is then aggregated to derive the nutritional data for the entire dish. Furthermore, in order to ensure a high-quality gold standard set for testing, we utilize human curation to isolate a superior subset specifically designated for testing purposes.
@@
Our proposed dataset facilitates large-scale training of LMMs for various food-related tasks within the same dataset. Nevertheless, this endeavor also presents new challenges associated with multi-task learning when fine-tuning large models. Mixture of Experts (MoE) (Huang et al. 2023; Dou et al. 2023; Liu et al. 2023b; Ponti et al. 2023) has been a common strategy in the field of Natural Language Processing (NLP) to handle multi-task fine-tuning. This method involves the utilization of multiple expert models, each specialized in different task or segment of the data distribution. Recently, there has been a surge in the application of MoE to LMMs. MoE-LLaVA (Lin et al. 2024), for instance, integrates MoE into its Feed-Forward Network (FFN) layers to enhance the model’s adaptability across tasks. On the other hand, LLaVA-MoLE (Chen, Jie, and Ma 2024) incorporates the MoE paradigm into the Low-Rank Adaptation (LoRA) (Hu et al. 2021) module, selecting the top-1 experts for each specific task to guarantee the spare activation of experts. However, as MoE-LLaVA incorporates experts throughout the FFN layer, the resulting increase in training parameters can be prohibitive. Conversely, although LLaVA-MoLE utilizes LoRA experts, which are more parameter-efficient, its strategy of selecting only the top expert may constrain the model’s flexibility on skill leverage. This is because certain tasks might share foundational skills. For example, in the food domain, both ingredients recognition and recipe generation are sensitive to the composition of ingredients in a dish.
@@
To address the aforementioned issues, we propose Linear Rectified Mixture of Diverse Experts (RoDE). Following LLaVA-MoLE (Chen, Jie, and Ma 2024), we design experts from the LoRA perspective to ensure efficient use of trainable parameters. Unlike LLaVA-MoLE, which dedicates each expert to a particular task, LLaVA-RoDE conceptualizes experts as granular skill modules, allowing multiple experts to participate in a single task. This design inspires us to develop experts with diverse capabilities—that is, we allocate varying amounts of trainable parameters to each expert in an effort to conserve GPU memory. To integrate the experts, we then introduce a linear rectified router to assign tasks to the appropriate LoRA experts. The router employs a Rectified Linear Unit (ReLU) to refine its output, which confers two significant benefits: 1) it enables sparse task-skill matching, which has been proved to surpass dense activation methods in effectiveness (Zhou et al. 2022; Chen, Jie, and Ma 2024); 2) the simplicity of the ReLU function makes it inherently easy to optimize. Consequently, this leads to a sparing activation of the diverse expert set. RoDE is engineered to optimize the allocation and usage of the model’s resources, significantly bolstering the capabilities of large-scale models in multi-task learning scenarios. Our empirical findings underscore the superior performance of RoDE in the realm of food multi-task learning challenges, validating the strength of our proposed methodology.
@@
Related Work
Large Multimodal Models (LMMs). In recent years, Large Language Models (LLMs) have showcased remarkable prowess in Natural Language Processing (NLP), paving the way for breakthroughs in multimodal learning. LMMs typically integrate a pre-trained vision encoder with a LLM architecture. Subsequently, visual features undergo adaptation via a projection module, facilitating their integration into the hidden space of LLMs for joint processing with textual inputs. Through multimodal training, LMMs acquire the capability to generate responses based on both visual and textual inputs. LLaVA (Liu et al. 2023a), for instance, introduces a vision encoder to LLMs and demonstrates significant enhancements across various vision-language tasks. Building upon LLaVA’s advancements, LISA (Lai et al. 2023) further enriches multimodal models by incorporating a segmentation module. This addition augments the model’s capacity to discern fine-grained details within visual inputs, resulting in more nuanced and contextually richer representations. GPT-4v (Achiam et al. 2023) stands out as one of the most powerful LMMs, capable of providing instructional data across numerous research domains. Its potency extends beyond NLP, contributing to advancements in various interdisciplinary fields.
Food Multi-task Learning. With food playing a significant part in human life, there has been a lot of work carried out doing research in this domain. In the early stage, food classification (Bossard, Guillaumin, and Van Gool 2014; Qiu et al. 2022) happened to be the most popular task related, bringing up a surge in the number of relevant datasets, like Food-101 (Bossard, Guillaumin, and Van Gool 2014), UEC Food 256 (Kawano and Yanai 2014) and Food2K (Min et al. 2023). As the tasks in the food domain developed into greater diversity, from ingredient recognition (Wang et al. 2022; Gao et al. 2023; Luo et al. 2023), recipe generation (Salvador et al. 2019; Chhikara et al. 2024) to nutrition estimation (Ando et al. 2019; Thames et al. 2021), food datasets have also grown to be more diverse and large-scale. Vireo Food-172 (Chen and Ngo 2016b) and 251 (Chen et al. 2021) are datasets containing not only category information but also ingredient labels, Recipe1M (Salvador et al. 2017) and Recipe1M+ (Marin et al. 2019) are datasets involving a wealth of recipe data, and Nutrition5K (Thames et al. 2021) is a dataset specialized in high accuracy nutritional content annotation. While all of the above methods have demonstrated their performance on food, they were all limited to address one or a few tasks, rather than integrating all tasks into a single multimodal model. To accomplish this, Yin et al. came up with FoodLMM (Yin et al. 2023), a versatile food assistant based on LMMs that could handle a variety of tasks in the food domain. However, their method might lead to task conflict when several tasks were fine-tuned at the same time, lacking the capability to efficiently implement multi-task learning.
@@
Table 1: Annotation Diversity in Food Datasets. The volume and variety of annotations differ significantly across various food datasets. Uni-Food offers a large-scale training dataset, accommodating a broad spectrum of food-related tasks.
Mixture of Experts (MoE) dynamically combines the decisions of multiple experts on the same input to improve overall performance. In LLMs, which typically adopt transformer architecture, MoE is often implemented within MLP layers. LoRAHub (Huang et al. 2023) initially trains a series of LoRA weights on upstream tasks. To adapt these to a downstream task, it employs a gradient-free method to search for the coefficients that will combine the pre-trained LoRA set. MOELoRA (Liu et al. 2023b), on the other hand, utilizes a router conditioned on a task identifier to dynamically merge multiple LoRA outputs. In contrast, MoCLE (Gou et al. 2023) designs a router that is conditioned on the clustering information of each individual input sample. LoRAMoE (Dou et al. 2023) partitions the LoRA experts into two groups and purposefully cultivates different capabilities within each group. All these mixture-of-LoRA methods have pre-set hyperparameters that require meticulous selection, and the LoRA experts are densely combined. (Zhou et al. 2022) conducted a comparison between the dense and sparse mixtures of LoRA experts for large language models, concluding that a dense mixture yields superior performance. Some studies have migrated MoE designs into LMMs. MoE-LLaVA (Lin et al. 2024) introduces multiple FFNs in stead of a single one and integrates a router mechanism to sample predictions. LLaVA-MoLE (Chen, Jie, and Ma 2024) applies multiple LoRAs and activates one specifically for a task. However, these methods fall short in effectively handling the nuances of rational, fine-grained task-skill allocation. Our paper aims to address this gap.
Dataset Construction
In this paper, we construct a large-scale dataset called Uni-Food. Different from other publicly available food datasets (Bossard, Guillaumin, and Van Gool 2014; Marin et al. 2019; Chen and Ngo 2016a; Thames et al. 2021), Uni-Food contains various attributes used in food-related tasks, including food category, ingredients, recipe and nutrition for each food image. To the best of our knowledge, this is the largest dataset that provides all the attributes in one dataset. Table 1 summarizes the tasks and sample sizes of existing primary datasets and our Uni-Food dataset.
Attribution
@@
Figure 3: Dataset Statistics.The category distribution of the dataset.
Figure 4: The illustration of our proposed Linear Rectified Mixture of Diverse Experts (RoDE) approach. We use LLaVA (Liu et al. 2023a) as the foundational LMM. RoDE module is incorporated into both the query projection layer and the value projection layer of each transformer block.
Nutrition Labeling
As the ingredient and recipe information can be easily collected from Recipe1M+ (Marin et al. 2019), we proceed to annotate the nutrition content. To acquire precise nutritional information, we feed both the food image and the ingredients with their respective amounts into ChatGPT4-vision  [^1] We intentionally omit recipe information to enhance processing efficiency, as we have observed that including recipes increases query time without considerably improving precision.. The output text is meticulously processed to extract the nutritional values. Leveraging the capabilities of ChatGPT4-vision, we collect detailed nutrition information at the ingredient level, encompassing metrics such as mass, fat content, protein content, carbohydrate content, and energy content. Importantly, the ingredient list in recipes contains the quantity for each ingredient, along with the food image, ChatGPT4-vision is capable of generating high-quality ingredient-level nutrition information according to our manual check based on USDA  [^2] https://fdc.nal.usda.gov/. The ingredient-level nutrition information is then aggregated to derive the comprehensive nutrition profile of the entire dish. Through this process, we obtain a holistic understanding of the nutritional composition of the dish, enabling precise estimation and analysis of its nutritional value.
Golden set selection
@@
Method
In this section, we first provide a preliminary overview of the problem formulation for multi-task tuning of LMMs, as well as the commonly used techniques, i.e., Low-Rank Adaptation (LoRA) (Hu et al. 2021) and Mixture of Experts (MoE) (Dou et al. 2023) for addressing this problem. Subsequently, we introduce our Linear Rectified Mixture of Diverse Experts approach in detail.
Preliminary
Problem Formulation.
A Large Multi-modal Model (LMM) can be constituted by a vision encoder and a Large Language Model (LLM), as shown in Figure 4. Normally, LMM is initially pre-trained with tremendous amounts of data, and then fine-tuned to adapt to downstream tasks. The utilization of multi-modal documents for Supervised Fine-Tuning (SFT) (Radford et al. 2018) is a common practice in the fine-tuning of LMMs.

### v3

@@ Abstract
Large Multi-modal Models (LMMs) have significantly advanced a variety of vision-language tasks. The scalability and availability of high-quality training data play a pivotal role in the success of LMMs. In the realm of food, while comprehensive food datasets such as Recipe1M offer an abundance of ingredient and recipe information, they often fall short of providing ample data for nutritional analysis. The Recipe1M+ dataset, despite offering a subset for nutritional evaluation, is limited in the scale and accuracy of nutrition information. To bridge this gap, we introduce Uni-Food, a unified food dataset that comprises over 100,000 images with various food labels, including categories, ingredients, recipes, and ingredient-level nutritional information. To mitigate the conflicts arising from multi-task supervision during fine-tuning of LMMs, we introduce a novel Linear Rectification Mixture of Diverse Experts (RoDE) approach. RoDE utilizes a diverse array of experts to address tasks of varying complexity, thereby facilitating the coordination of trainable parameters, i.e., it allocates more parameters for more complex tasks and, conversely, fewer parameters for simpler tasks. RoDE implements linear rectification union to refine the router’s functionality, thereby enhancing the efficiency of sparse task allocation. These design choices endow RoDE with features that ensure GPU memory efficiency and ease of optimization. Our experimental results validate the effectiveness of our proposed approach in addressing the inherent challenges of food-related multitasking. Project page at: https://pengkun-jiao.github.io/UniFood-project/.
@@
Figure 2: Composition of Labels in the UniFood Dataset. The annotation consists of category, ingredient list, recipe instructions, and nutrition information.
I Introduction
@@
Food occupies a central position in our daily lives, leading to the emergence of various food-related tasks, e.g., ingredient recognition, recipe generation, and nutritional estimation [1]. These tasks have attracted considerable research interests over the years [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Building on the success of Large Language Models (LLMs) [15, 16], Large Multi-modal Models [17, 10] have begun to make a significant impact in many specialized areas, including the food domain [18].
@@
The success of LLMs and LMMs can be largely attributed to the availability of large-scale training data. In the food domain, while Recipe1M [19] provides over a million web-crawled recipes and thousands of food images, it lacks comprehensive nutritional information. While Recipe1M+ [20] includes a subset containing nutritional information, it suffers from limitations in data scale and annotation quality. Similarly, nutrition-focused datasets such as Nutrition5k [21] are also limited in scale and may suffer from domain shift issues, as the labeled data primarily consists of lightly processed ingredients rather than fully cooked meals. Moreover, previous studies [22, 23] have indicated that integrating diverse sources of data for training can lead to data conflicts. The performance of an LMM on a specific task also depends heavily on the representation of that task’s data within the overall training dataset [23]. These challenges highlight the need for a unified dataset covering all food-related tasks from the same source, aiming at mitigating data conflicts and ensuring balanced representation across tasks.
@@
In order to overcome these obstacles, we introduce Uni-Food, a unified dataset that encapsulates various food-related information, i.e., food category, ingredients, recipes, and valuable ingredient-level nutritional information for each food image. Specifically, we curate 100,000 samples from Recipe1M [19] and employ ChatGPT-4 [24] to generate nutritional information for each ingredient list. This information is then aggregated to derive the nutritional data for the entire dish. Furthermore, in order to ensure a high-quality gold standard set for testing, we utilize human curation to isolate a superior subset specifically designated for testing purposes.
@@
Our proposed dataset facilitates large-scale training of LMMs for various food-related tasks within the same dataset. Nevertheless, this endeavor also presents new challenges associated with multi-task learning when fine-tuning large models. Mixture of Experts (MoE) [25, 26, 27, 28] has been a common strategy in the field of Natural Language Processing (NLP) to handle multi-task fine-tuning. This method involves the utilization of multiple expert models, each specialized in different task or segment of the data distribution. Recently, there has been a surge in the application of MoE to LMMs. MoE-LLaVA [29], for instance, integrates MoE into its Feed-Forward Network (FFN) layers to enhance the model’s adaptability across tasks. On the other hand, LLaVA-MoLE [23] incorporates the MoE paradigm into the Low-Rank Adaptation (LoRA) [30] module, selecting the top-1 experts for each specific task to guarantee the spare activation of experts. However, as MoE-LLaVA incorporates experts throughout the FFN layer, the resulting increase in training parameters can be prohibitive. Conversely, although LLaVA-MoLE utilizes LoRA experts, which are more parameter-efficient, its strategy of selecting only the top expert may constrain the model’s flexibility on skill leverage. This is because certain tasks might share foundational skills. For example, in the food domain, both ingredients recognition and recipe generation are sensitive to the composition of ingredients in a dish.
@@
To address the aforementioned issues, we propose Linear Rectified Mixture of Diverse Experts (RoDE). Following LLaVA-MoLE [23], we design experts from the LoRA perspective to ensure efficient use of trainable parameters. Unlike LLaVA-MoLE, which dedicates each expert to a particular task, LLaVA-RoDE conceptualizes experts as granular skill modules, allowing multiple experts to participate in a single task. This design inspires us to develop experts with diverse capabilities—that is, we allocate varying amounts of trainable parameters to each expert in an effort to conserve GPU memory. To integrate the experts, we then introduce a linear rectified router to assign tasks to the appropriate LoRA experts. The router employs a Rectified Linear Unit (ReLU) to refine its output, which confers two significant benefits: 1) it enables sparse task-skill matching, which has been proved to surpass dense activation methods in effectiveness [22, 23]; 2) the simplicity of the ReLU function makes it inherently easy to optimize. Consequently, this leads to a sparing activation of the diverse expert set. RoDE is engineered to optimize the allocation and usage of the model’s resources, significantly bolstering the capabilities of large-scale models in multi-task learning scenarios. Our empirical findings underscore the superior performance of RoDE in the realm of food multi-task learning challenges, validating the strength of our proposed methodology.
@@
II Related Work
Large Multimodal Models (LMMs). In recent years, Large Language Models (LLMs) have showcased remarkable prowess in Natural Language Processing (NLP), paving the way for breakthroughs in multimodal learning. LMMs typically integrate a pre-trained vision encoder with a LLM architecture. Subsequently, visual features undergo adaptation via a projection module, facilitating their integration into the hidden space of LLMs for joint processing with textual inputs. Through multimodal training, LMMs acquire the capability to generate responses based on both visual and textual inputs. LLaVA [17], for instance, introduces a vision encoder to LLMs and demonstrates significant enhancements across various vision-language tasks. Building upon LLaVA’s advancements, LISA [31] further enriches multimodal models by incorporating a segmentation module. This addition augments the model’s capacity to discern fine-grained details within visual inputs, resulting in more nuanced and contextually richer representations. GPT-4v [24] stands out as one of the most powerful LMMs, capable of providing instructional data across numerous research domains. Its potency extends beyond NLP, contributing to advancements in various interdisciplinary fields.
Food Multi-task Learning. With food playing a significant part in human life, there has been a lot of work carried out doing research in this domain. In the early stage, food classification [32, 33] happened to be the most popular task related, bringing up a surge in the number of relevant datasets, like Food-101 [32], UEC Food 256 [34] and Food2K [35]. As the tasks in the food domain developed into greater diversity, from ingredient recognition [9, 36, 10], recipe generation [37, 38] to nutrition estimation [39, 21], food datasets have also grown to be more diverse and large-scale. Vireo Food-172 [40] and 251 [6] are datasets containing not only category information but also ingredient labels, Recipe1M [19] and Recipe1M+ [20] are datasets involving a wealth of recipe data, and Nutrition5K [21] is a dataset specialized in high accuracy nutritional content annotation. While all of the above methods have demonstrated their performance on food, they were all limited to address one or a few tasks, rather than integrating all tasks into a single multimodal model. To accomplish this, Yin et al. came up with FoodLMM [18], a versatile food assistant based on LMMs that could handle a variety of tasks in the food domain. However, their method might lead to task conflict when several tasks were fine-tuned at the same time, lacking the capability to efficiently implement multi-task learning.
TABLE I: Annotation Diversity in Food Datasets. The volume and variety of annotations differ significantly across various food datasets. Uni-Food offers a large-scale training dataset, accommodating a broad spectrum of food-related tasks.
@@ II Related Work
Mixture of Experts (MoE) dynamically combines the decisions of multiple experts on the same input to improve overall performance. In LLMs, which typically adopt transformer architecture, MoE is often implemented within MLP layers. LoRAHub [25] initially trains a series of LoRA weights on upstream tasks. To adapt these to a downstream task, it employs a gradient-free method to search for the coefficients that will combine the pre-trained LoRA set. MOELoRA [27], on the other hand, utilizes a router conditioned on a task identifier to dynamically merge multiple LoRA outputs. In contrast, MoCLE [41] designs a router that is conditioned on the clustering information of each individual input sample. LoRAMoE [26] partitions the LoRA experts into two groups and purposefully cultivates different capabilities within each group. All these mixture-of-LoRA methods have pre-set hyperparameters that require meticulous selection, and the LoRA experts are densely combined. [22] conducted a comparison between the dense and sparse mixtures of LoRA experts for large language models, concluding that a dense mixture yields superior performance. Some studies have migrated MoE designs into LMMs. MoE-LLaVA [29] introduces multiple FFNs in stead of a single one and integrates a router mechanism to sample predictions. LLaVA-MoLE [23] applies multiple LoRAs and activates one specifically for a task. However, these methods fall short in effectively handling the nuances of rational, fine-grained task-skill allocation. Our paper aims to address this gap.
Figure 3: The nutritional statistics (per 100g) and ingredient statistics for UniFood.
III Dataset Construction
In this paper, we construct a large-scale dataset called Uni-Food. Different from other publicly available food datasets [32, 20, 40, 21], Uni-Food contains various attributes used in food-related tasks, including food category, ingredients, recipe and nutrition for each food image. To the best of our knowledge, this is the largest dataset that provides all the attributes in one dataset. Table I summarizes the tasks and sample sizes of existing primary datasets and our Uni-Food dataset.
III-A Attribution
@@
Figure 4: UniFood Dataset Statistics. The category distribution of the dataset.
Figure 5: The illustration of our proposed Linear Rectified Mixture of Diverse Experts (RoDE) approach. We use LLaVA [17] as the foundational LMM. RoDE module is incorporated into both the query projection layer and the value projection layer of each transformer block.
III-B Nutrition Labeling
As the ingredient and recipe information can be easily collected from Recipe1M+ [20], we proceed to annotate the nutrition content. To acquire precise nutritional information, we feed both the food image and the ingredients with their respective amounts into ChatGPT4-vision  [^1] We intentionally omit recipe information to enhance processing efficiency, as we have observed that including recipes increases query time without considerably improving precision.. The output text is meticulously processed to extract the nutritional values. Leveraging the capabilities of ChatGPT4-vision, we collect detailed nutrition information at the ingredient level, encompassing metrics such as mass, fat content, protein content, carbohydrate content, and energy content. Importantly, the ingredient list in recipes contains the quantity for each ingredient, along with the food image, ChatGPT4-vision is capable of generating high-quality ingredient-level nutrition information according to our manual check based on USDA  [^2] https://fdc.nal.usda.gov/. For detailed prompt design, please refer to the Sec. A of the supplementary materials. Subsequently, the ingredient-level nutrition information is then aggregated to derive the comprehensive nutrition profile of the entire dish. Through this process, we obtain a holistic understanding of the nutritional composition of the dish, enabling precise estimation and analysis of its nutritional value.
III-C Golden set selection
@@
III-D Nutrition annotation quality
We ensured high-quality nutrition annotation in the ChatGPT-4V by supplying it with both the quantity and name of each ingredient from the recipe, as verified through our earlier manual checks. To evaluate the disparity between ChatGPT annotation and human annotation, we randomly select 1,000 entries from our golden nutrition test set and prompt ChatGPT-4V to generate their nutrition information. The results, detailed in Table II, show minimal discrepancies between the human-verified values and those generated by ChatGPT-4V, demonstrating the high quality of our annotations. For further insights, case studies on nutritional annotation are discussed in Sec. B of the supplementary material.
TABLE II: Gap between Human and ChatGPT-4V Annotations.
Metric Mass (g) Calories (kcal) Fat (g) Protein Value (g)
MAE 13.22 54.42 2.67 2.39
pMAE 1.77% 2.7% 2.59% 6.91%
IV Method
In this section, we first provide a preliminary overview of the problem formulation for multi-task tuning of LMMs, as well as the commonly used techniques, i.e., Low-Rank Adaptation (LoRA) [30] and Mixture of Experts (MoE) [26] for addressing this problem. Subsequently, we introduce our Linear Rectified Mixture of Diverse Experts approach in detail.
IV-A Preliminary
IV-A1 Problem Formulation.
A Large Multi-modal Model (LMM) can be constituted by a vision encoder and a Large Language Model (LLM), as shown in Figure 5. Normally, LMM is initially pre-trained with tremendous amounts of data, and then fine-tuned to adapt to downstream tasks. The utilization of multi-modal documents for Supervised Fine-Tuning (SFT) [15] is a common practice in the fine-tuning of LMMs.